OpenStudy (anonymous):

how do you find soultions for 30x^2-28x+6=0

OpenStudy (dogzcatz):

is it Multiple choice or neah

OpenStudy (anonymous):

yeah

OpenStudy (dogzcatz):

What are the choices

OpenStudy (anonymous):

x=2/3 x=1/5 x=3/5 x=1/2 x=1/3 x=4/5

OpenStudy (dogzcatz):

3/5 and 1/3 are the answers

OpenStudy (anonymous):

thanks

OpenStudy (dogzcatz):

I had those answers before, but just wanted to know

OpenStudy (anonymous):

ah okay

OpenStudy (anonymous):

what about x^2-9=0

OpenStudy (anonymous):

do ur work lol

OpenStudy (anonymous):

would it be 3 and -3

OpenStudy (mathstudent55):

To find solutions to this quadratic equation, you can factor the polynomial, or you can use the quadratic formula.